# Extrinsic Evaluation of Cultural Competence in Large Language Models

### v2

@@ Abstract
Shaily Bhatt Carnegie Mellon University shaily@cmu.edu                        Fernando Diaz Carnegie Mellon University diazf@cmu.edu
@@ 1 Introduction
Cultural competence is the ability to effectively and appropriately communicate with socioculturally different audiences Deardorff (2009).111interchangeably, the terms intercultural competence and crosscultural competence are also used. People demonstrate cultural competence by tailoring their utterances to the participants in a conversation Bell (1984); Hawkins et al. (2021); Wu et al. (2023a). These adaptations range from sociolinguistic variations (e.g., using ‘soccer’ or ‘football’ depending on the context) to appropriately using facts (e.g., in India, the Prime Minister is the head of the government, but in USA, the President is). Hence, for effectively serving diverse users, outputs from large language models (LLMs) need to be culturally relevant Hovy and Yang (2021).
Cultural competence consists of multiple components, including a person’s knowledge of a culture, which then supplements their skills of effectively communicating with people from that culture Deardorff (2006); Fantini and Tirmizi (2006); Alizadeh and Chavan (2016). So, cultural competence of LLMs should also be evaluated along both these aspects. Contemporary works have largely targeted the knowledge component of cultural competence by evaluating LLMs’ knowledge of cultural values, norms, and artefacts (§ 2.2). Such evaluation is intrinsic because it is decoupled from the manifestation of this knowledge in downstream applications Jones and Galliers (1995).
In this work, we focus on extrinsic evaluation of cultural competence. Extrinsic evaluation setups should closely mimic user interactions with a system Jones and Galliers (1995). We select the tasks of story generation and open-ended question answering (QA), both of which have high representation in user interactions with LLMs Zhao et al. (2024). We evaluate the lexical variations in outputs of 6 LLMs for 195 nationalities, and by proxy culture, for these tasks using both qualitative and quantitative analyses. Further, recent intrinsic evaluations have heavily relied on surveys from crosscultural psychology, like Hofstede’s Cultural Dimensions Hofstede et al. (2010) and World Values Survey Haerpfer et al. (2022), as a measure of cultural values across countries Arora et al. (2023); Cao et al. (2023); Durmus et al. (2023); Ramezani and Xu (2023); AlKhamissi et al. (2024); Masoud et al. (2024). Thus, we evaluate whether the text distributions of outputs correlate with the cultural values of countries, as captured by these surveys. Our three main research questions are:
@@ 3. RQ3:
By measuring the variance in the outputs, we find that models make non-trivial adaptations for different nationalities (§ 5.1). Next, inspecting the vocabulary of these outputs, we find that they contain culturally relevant words (§ 5.2). Finally, we find only a weak correlation between the text distributions and cultural values of countries, as measured by crosscultural psychology surveys frequently used in contemporary work (§ 5.3).
@@
All our code and data will be open-sourced.
@@ 2.1 Cultural Competence
Cultural competence is the ability to effectively communicate with a socioculturally different audience Deardorff (2009). While multiple definitions exist Alizadeh and Chavan (2016), agreed-upon components include 1 the awareness about one’s positionality and attitude, 2 the knowledge about the language, values, beliefs, practices, symbols etc. of a culture, and 3 the skill of appropriately using this knowledge when communicating Howard-Hamilton et al. (1998); Deardorff (2006); Fantini and Tirmizi (2006); Deardorff (2009). [^2] For LLMs, we only rely on analogy to ‘knowledge’ and ‘skills’, and do not invoke analogies to ‘awareness’.
The knowledge component requires understanding differences in values, beliefs, and preferences across societies. Surveys in crosscultural psychology, like Hofstede’s Cultural Dimensions (HCD) Hofstede (2001) and World Values Survey (WVS) Haerpfer et al. (2022) attempt to elicit these differences across cultures, proxied by nationalities, using value-based questions. [^3] For example one of the questions in the Hofstede’s survey is ”In choosing an ideal job, how important would it be to you to have sufficient time for your personal or home life?”. Survey responses from a large number of individuals are used to quantify the differences in cultural values across countries. Hofstede’s theory, in particular, has been widely adopted in fields requiring cultural competence such as communication, education, business, and healthcare Ahern et al. (2012); Burai (2016); Chang and Wu (2023); Singh and Kumari (2023).  [^4] We note that defining the underpinning concept of culture itself remains elusive. Numerous works have attempted to synthesize the definitions of culture across disciplines, highlighting its complex and multi-faceted nature Kroeber and Kluckhohn (1952); Baldwin et al. (2006). Broadly, culture is a shared collection of knowledge, values, practices, norms, and beliefs that manifest in expression as behavioural and linguistic patterns Kroeber and Kluckhohn (1952).
@@ 2.2 Cultural Competence in LLMs
There is a growing body of work on ensuring that LLMs align with diverse human values Hershcovich et al. (2022); Wu et al. (2023b); Kirk et al. (2024); Sorensen et al. (2024) and can serve socioculturally diverse users Hovy and Yang (2021); Hershcovich et al. (2022); Adilazuarda et al. (2024). Specifically, prior works have evaluated LLMs for:
1. Reflection of diverse cultural values on crosscultural psychology surveys (like HCD and WVS) using MCQs, Chain of Thought prompting, or personas Arora et al. (2023); Cao et al. (2023); Durmus et al. (2023); Ramezani and Xu (2023); AlKhamissi et al. (2024); Masoud et al. (2024).
@@
3. Commonsense and figurative language understanding using MCQs Nguyen et al. (2023); Palta and Rudinger (2023); Kabra et al. (2023); Kim et al. (2024); Koto et al. (2024); Wang et al. (2024), and pragmatic games Shaikh et al. (2023).
@@
These works reveal gaps in LLMs’ knowledge of non-western cultures, complimenting known demographic biases in LLMs Mishra et al. (2020); Zhou et al. (2022); Basu et al. (2023); Jha et al. (2023); Schwöbel et al. (2023); Naous et al. (2024).
@@ 3 Extrinsic Evaluation of Cultural Competence
Jones and Galliers (1995) describe extrinsic evaluation criteria as, ‘those relating to its function, i.e its role in relation to its setup’s purpose’. So, consider the two broad use cases of LLMs: 1 classification tasks, and 2 generation tasks. While incorporating cultural knowledge has been shown to benefit classification tasks like hate-speech detection and commonsense reasoning Zhou et al. (2023); Li et al. (2024a); Shi et al. (2024), to the best of our knowledge there is no prior work focusing on open-ended text generation tasks.
@@ 3.1 Quantitative Evaluation
In order to quantify how much the generated language varies when nationalities are perturbed, we measure the variance in distance between outputs, where distance is computed according to a specific representation (§ 4.3.1).
@@ 4.1 Tasks and Data
We select two tasks, story generation and open-ended question answering for our experiments. These were selected as they fulfil two main criteria. First, they have a sizeable representation in user interactions with LLMs Zhao et al. (2024). And second, they represent diverse types of generation tasks with story generation on the creative end of the spectrum, while open-ended question answering being on the factual end of the spectrum.
@@
We created a list of 345 topics across 13 categories. We selected the categories (biology, chemistry, economics, environment, humanities, history, law, maths, physics, politics, religion, space, and world affairs) to ensure diversity in topics. Next, we curated topics for each category by referring to textbooks and encyclopedias. [^5] The datasheet Gebru et al. (2021) is in Appendix A and curation logs will be released with the data upon publication. Examples of topics include: ‘elections’ in ‘politics’, ‘inertia’ in ‘physics’, ‘photosynthesis’ in ‘biology’. For this task, use a simple prompt template:
@@
These results in prompts like ‘Explain elections to an Indian Person in English’. [^6] We observed that not including the phrase ‘in English’ in the prompt resulted in GPT 3.5’s output often being in the dominant language of the particular country, for example for ‘Mexican’ the output is in Spanish. While this is an interesting phenomenon, analyzing this is beyond the scope of this paper.
@@
We created a list of 35 topics for children’s stories. We used online websites and children’s storybooks to come up with topics. Examples include topics like moral values (‘honesty’, ‘kindness’), characters (‘farm animals’, ‘birds’), and places (‘school’, ‘jungle’). Similar to QA, we use a simple prompt template:
@@ 4.2 Models
We evaluate the following LLMs: 1 GPT 3.5 Turbo (gpt-3.5-turbo-0125) [^7] https://platform.openai.com/docs/models/gpt-3-5-turbo, queried via API between February 23 and March 28 2024. 2 Gemma 2B instruct and 7B instruct Team et al. (2024) 3 Llama 2 7B chat and 13B chat Touvron et al. (2023) 4 Llama 3 8B instruct AI@Meta (2024) . We sample 5 responses per prompt, using a temperature of 0.3. We generate a maximum of 100 tokens for QA and 1000 tokens for stories.
@@ 4.3.1 Text Similarity
We picked BLEU and WED to focus on capturing the differences in lexical items between two outputs, e.g., the use of ‘soccer’ or ‘football’. [^8] In early expeirments we found that semantic metrics like BERTscore Zhang* et al. (2020) or embedding similarity might not be suitable because: 1 a lot of culturally relevant words from the outputs were converted to [UNK] tokens, 2 we did not see differences in the embedding for outputs that were qualitatively different, especially in QA; perhaps partly because of (a) and and because, intuitively the the different words convey the same meaning.
@@ 4.3.2 Difference in Cultural Values
Following prior work, we rely on data from crosscultural psychology surveys to measure the difference in cultural values among countries.
@@
We use data from 64 countries and represent each country with 249 dimensions using the 249 questions from WVS [^10] There are additional questions that are either non-ordinal or descriptive in nature or are experimental, which we ignore. [^11] https://www.worldvaluessurvey.org/
WVSDocumentationWV7.jsp
@@
Chinese li , mei , china , ming , chen , wu , wei , xiao , wukong , feather ,
ping , lake , bao , snowball , chinese
German hans , germany , lena , anna , fritz , max , gretchen , bauer , lorelei ,
herr , lila , liesl , rübezahl , emma , karl
Indian raj , india , rani , arjun , ravi , priya , guru , peacock , krishna ,
raja , meena , gupta , durga , beggar , temple
Nigerian kola , nigeria , tunde , bola , kemi , ade , oya , adaeze , ayo ,
zuri , lagos , jide , nigerian , simba , heron
@@
Figure 2: Kendall’s $\tau_{c}$ rank correlation between text distribution and cultural closeness of countries. For both plots, text similarity is measured using BLEU. For HCD correlation statistic values are greater than 0, implying a small but positive correlation (2(a)). However, for WVS, most correlations are less than 0, indicating small and negative correlation (2(b)). There are no clear trends among different models or tasks.
@@ 5.1 Variance due to Nationality Perturbation
Our first research question was to analyse the extent of variation in outputs when nationalities are perturbed in the prompt. For this, we quantify the lexical variance (§ 3.1) in outputs, as measured by word edit distance in Figure 1. We find that model outputs do vary with changing nationalities for both tasks across models. Moreover, these variations are non-trivial and task dependent, as described below.
Control experiment: variance within nationality
We want to ensure that that the variance observed across nationalities are non-trivial, i.e. they do not occur because of the non-deterministic nature of generation in models. For this, we also measure the variance within multiple outputs for a particular nationality. We find that the variance for outputs within nationality is consistently lower than the variance across nationalities. We confirm this with ANOVA having a p-value of <0.05 (Appendix B.2).
@@
We find that the nature of the task affects the extent of variation. The median variance for story generation is higher than the median variance for QA for every model. This might be expected as story generation, had longer outputs and being a creative task allows for more adaptations. On the other hand, the difference between the upper and lower quartiles of variance for QA is larger than that for stories. This is likely because QA consists of a wider variety of topics ranging from scientific categories, where limited variations might be expected, to topics on politics and history, that allow more variation in answers than others. For example answers while explaining ‘elections’ (politics) might vary more as they are operationalized differently across countries, but explaining ‘inertia’ (physics) might not vary as much.
@@ 5.2 Culturally Relevant Words in Outputs
For the topics in the politics category of the QA task, we see words referring to senate houses and political offices of the countries, for example, ‘lok sabha’ and ‘rajya sabha’ in India, ‘bundestag’ for Germany, and ‘meshrano jirga’ and ‘wolesi jirga’ for Afghanistan. The list also features politically polarised issues such as ‘gun’ in America and ‘brexit’ in UK. Another common feature is the names of political parties, such as ‘bjp’ in India, ‘apc’ and ‘pdp’ in Nigeria, and ‘ndp’ in Canada.
Finally, we note that the cultural relevance of all the words on the lists is not obvious (e.g ‘notably’ in German in Table 2). Moreover, not all topics in the QA setting surface such interpretable lists of culturally relevant words. Especially lexicon from scientific topics did not reveal interesting examples when inspecting the top-scoring TF-IDF words. This further compliments our earlier finding of output variations being different across tasks.

### v3

@@ Abstract
Shaily Bhatt Carnegie Mellon University shaily@cmu.edu                        Fernando Diaz Carnegie Mellon University diazf@acm.org
@@ 1 Introduction
Figure 1: Outputs for Question Answering and Story Generation vary when explicit cue of culture, i.e. nationality, in the prompt is perturbed. We collect outputs for 345 topics for QA, 35 topics for stories, 193 nationalities, 6 LLMs, 5 responses per prompt, and 2 temperatures. We then evaluate these outputs for the extent of lexical variance (§5.1), culturally relevant vocabulary (§5.2), and correlation between text distribution and the cultural values (§5.3).
Cultural competence is the ability to effectively and appropriately communicate with socioculturally different audiences Deardorff (2009). People demonstrate cultural competence by tailoring their utterances to the participants in a conversation Bell (1984); Hawkins et al. (2021); Wu et al. (2023a). These adaptations range from sociolinguistic variations (e.g., using ‘soccer’ or ‘football’ depending on the context) to appropriately using facts (e.g., in India, the Prime Minister is the head of the government, but in the USA, the President is). Hence, for effectively serving diverse users, outputs from large language models (LLMs) need to be culturally relevant Hovy and Yang (2021).
Cultural competence includes a person’s knowledge of a culture, which then supplements their skills of effectively communicating with people from that culture Deardorff (2006); Fantini and Tirmizi (2006); Alizadeh and Chavan (2016). So, the cultural competence of LLMs should also be evaluated along both these aspects. Contemporary works have largely targeted the knowledge component of cultural competence by evaluating LLMs’ knowledge of cultural values, norms, and artefacts (§ 2.2). Such evaluation is intrinsic because it is decoupled from the manifestation of this knowledge in downstream applications Jones and Galliers (1995).
In this work, we focus on extrinsic evaluation of cultural competence. Extrinsic evaluation setups should closely mimic user interactions with a system Jones and Galliers (1995). We select the tasks of story generation and open-ended question answering (QA), both of which have high representation in user interactions with LLMs Zhao et al. (2024), so . As shown in Figure 1, we obtain model outputs for 35 topics in story generation and 345 topics for QA when an explicit cue of culture, i.e., nationality, is present in the prompt. Our data consists of outputs from 6 LLMs for 193 nationalities, with 5 outputs per prompt and 2 temperature settings, resulting in over 370K story outputs and 3.6M QA outputs. We analyse the extent and the nature of lexical variations in these outputs. Further, recent intrinsic evaluations have heavily relied on surveys from cross-cultural psychology, like Hofstede’s Cultural Dimensions Hofstede et al. (2010) and World Values Survey Haerpfer et al. (2022) as measures of cultural values (e.g., Arora et al. (2023); DURMUS et al. (2024); AlKhamissi et al. (2024), inter alia). We evaluate whether the text distributions of outputs correlate with the cultural values of countries, as captured by these surveys. To summarize, our main research questions are:
@@ 3. RQ3:
By measuring the variance in the outputs, we find that models make non-trivial adaptations for different nationalities (§ 5.1). Next, inspecting the vocabulary of these outputs, we find that they contain culturally relevant words (§ 5.2). Finally, we find only a weak correlation between the text distributions and cultural values of countries, as measured by cross-cultural psychology surveys frequently used in contemporary work (§ 5.3).
@@
We make our code [^1] https://github.com/shaily99/eecc and data [^2] https://huggingface.co/datasets/shaily99/eecc publicly available for replicating results and use in further analysis.
@@ 2.1 Cultural Competence
Cultural competence333interchangeably, the terms intercultural competence and cross-cultural competence are also used. is the ability to effectively communicate with a socioculturally different audience Deardorff (2009). While multiple definitions exist Alizadeh and Chavan (2016), agreed-upon components include 1 the awareness about one’s positionality and attitude, 2 the knowledge about the language, values, beliefs, practices, symbols, etc. of a culture, and 3 the skill of appropriately using this knowledge when communicating Howard-Hamilton et al. (1998); Deardorff (2006); Fantini and Tirmizi (2006); Deardorff (2009). [^4] For LLMs, we only rely on analogy to ‘knowledge’ and ‘skills’, and do not invoke analogies to ‘awareness’.
The knowledge component requires understanding differences in values, beliefs, and preferences across societies. Surveys in cross-cultural psychology, like Hofstede’s Cultural Dimensions (HCD) Hofstede (2001) and World Values Survey (WVS) Haerpfer et al. (2022) attempt to elicit these differences across cultures, proxied by nationalities, using value-based questions. [^5] For example one of the questions in the Hofstede’s survey is “In choosing an ideal job, how important would it be to you to have sufficient time for your personal or home life?”. Survey responses from a large number of individuals are used to quantify the differences in cultural values across countries. Hofstede’s theory, in particular, has been widely adopted in fields requiring cultural competence such as communication, education, business, and healthcare Ahern et al. (2012); Burai (2016); Chang and Wu (2023); Singh and Kumari (2023).  [^6] We note that defining the underpinning concept of culture itself remains elusive. Numerous works have attempted to synthesize the definitions of culture across disciplines, highlighting its complex and multi-faceted nature Kroeber and Kluckhohn (1952); Baldwin et al. (2006). Broadly, culture is a shared collection of knowledge, values, practices, norms, and beliefs that manifest in expression as behavioural and linguistic patterns Kroeber and Kluckhohn (1952).
@@ 2.2 Cultural Competence in LLMs
There is a growing body of work on ensuring that LLMs align with diverse human values Hershcovich et al. (2022); Wu et al. (2023b); Kirk et al. (2024); Sorensen et al. and can serve socioculturally diverse users Hovy and Yang (2021); Hershcovich et al. (2022); Adilazuarda et al. (2024). Specifically, prior works have evaluated LLMs for:
1. Reflection of diverse cultural values on cross-cultural psychology surveys (like HCD and WVS) using MCQs, Chain of Thought prompting, or personas Arora et al. (2023); Cao et al. (2023); DURMUS et al. (2024); Ramezani and Xu (2023); AlKhamissi et al. (2024); Masoud et al. (2024).
@@
3. Commonsense and figurative language understanding using MCQs Palta and Rudinger (2023); Kabra et al. (2023); Kim et al. (2024); Koto et al. (2024); Wang et al. (2024), and pragmatic games Shaikh et al. (2023).
@@
These works reveal gaps in LLMs’ knowledge of non-western cultures, complementing known demographic biases in LLMs Mishra et al. (2020); Zhou et al. (2022); Basu et al. (2023); Jha et al. (2023); Schwöbel et al. (2023); Naous et al. (2024).
@@ 3 Extrinsic Evaluation of Cultural Competence
Jones and Galliers (1995) describe extrinsic evaluation criteria as, “those relating to its function, i.e its role in relation to its setup’s purpose”. So, consider the two broad use cases of LLMs: 1 classification tasks, and 2 generation tasks. While incorporating cultural knowledge has been shown to benefit classification tasks like hate speech detection and commonsense reasoning Zhou et al. (2023); Li et al. (2024a); Shi et al. (2024), to the best of our knowledge there is no prior work focusing on open-ended text generation tasks.
@@ 3.1 Quantitative Evaluation
In order to quantify how much the generated language varies when nationalities are perturbed, we measure the variance in word edit distance between outputs.
@@ 4.1 Tasks and Data
We select two tasks, story generation and open-ended question answering for our experiments. These were selected as they fulfil two main criteria. First, they have a sizeable representation in user interactions with LLMs Zhao et al. (2024). Second, they represent diverse types of generation tasks with story generation on the creative end of the spectrum, while question answering being on the factual end of the spectrum.
@@
We created a list of 345 topics across 13 categories. We selected the categories (biology, chemistry, economics, environment, humanities, history, law, maths, physics, politics, religion, space, and world affairs) to ensure diversity in topics. Next, we curated topics for each category by referring to 1 textbooks used for ACT and SAT Prep in the US, and NCERT textbooks that are used commonly in high schools in India; and 2 index terms of “topics” of various categories that were available on Wikipedia and Britannica.  [^7] The datasheet Gebru et al. (2021) is in Appendix D. Examples of topics include: ‘elections’ in ‘politics’, ‘inertia’ in ‘physics’, ‘photosynthesis’ in ‘biology’. For this task, use a simple prompt template:
@@
These results in prompts like ‘Explain elections to an Indian person in English.’ [^8] We observed that not including the phrase ‘in English’ in the prompt resulted in GPT 3.5’s output often being in the dominant language of the particular country, for example for ‘Mexican’ the output is in Spanish. While this is an interesting phenomenon, analyzing this is beyond the scope of this paper.
@@
We created a list of 35 topics for children’s stories. We used online websites of to curate the list, specifically, freechildrenstories.com, storyberries.com, and parenteducate.com. Examples include topics like moral values (‘honesty’, ‘kindness’), characters (‘farm animals’, ‘birds’), and places (‘school’, ‘jungle’). Similar to QA, we use a simple prompt template:
@@ 4.2 Models
We evaluate the following LLMs: 1 GPT 3.5 Turbo (gpt-3.5-turbo-0125), [^9] https://platform.openai.com/docs/models/gpt-3-5-turbo queried via API between February 23 and March 28 2024. 2 Gemma 2B instruct and 7B instruct Team et al. (2024) 3 Llama 2 7B chat and 13B chat Touvron et al. (2023) 4 Llama 3 8B instruct AI@Meta (2024) . We sample 5 responses per prompt, with a temperature of 0.3 and 0.7 (for all except GPT 3.5). We generate a maximum of 100 tokens for QA outputs and 1000 tokens for story outputs.
@@ 4.3.1 Text Similarity
We picked BLEU and WED to focus on capturing the differences in lexical items between two outputs, e.g., the use of ‘soccer’ or ‘football’. [^10] In early experiments we found that semantic metrics like BERTscore Zhang et al. (2020) or embedding similarity might not be suitable because: 1 a lot of culturally relevant words from the outputs were converted to [UNK] tokens, 2 we did not see differences in the embedding for outputs that were qualitatively different, especially in QA; perhaps partly because of (a) and because, intuitively the the different words convey the same meaning.
@@ 4.3.2 Difference in Cultural Values
Following prior work, we rely on data from cross-cultural psychology surveys to measure the differences in cultural values among countries.
@@
We use data from 64 countries and represent each country with 249 dimensions using the 249 questions from WVS. [^12] There are additional questions that are either non-ordinal or descriptive in nature or are experimental, which we ignore. [^13] https://www.worldvaluessurvey.org/WVSDocumentationWV7.jsp
@@
Chinese li, mei, china, ming, chen, wu, wei, xiao, wukong, feather, ping, lake, bao, snowball, chinese
German hans, germany, lena, anna, fritz, max, gretchen, bauer, lorelei, herr, lila, liesl, rübezahl, emma, karl
Indian raj, india, rani, arjun, ravi, priya, guru, peacock, krishna, raja, meena, gupta, durga, beggar, temple
Nigerian kola, nigeria, tunde, bola, kemi, ade, oya, adaeze, ayo, zuri, lagos, jide, nigerian, simba, heron
@@
Figure 3: Kendall’s $\tau_{c}$ rank correlation between text distribution and cultural closeness of countries. For both plots, text similarity is measured using BLEU. For HCD correlation statistic values are greater than 0, implying a small but positive correlation (6(a)). However, for WVS, most correlations are less than 0, indicating small and negative correlation (6(b)). There are no clear trends among different models or tasks.
@@ 5 Results
We now describe our observations on the outputs obtained with the temperature of 0.3. The observations with a temperature of 0.7 are consistent with these findings and are described in appendix C.
@@ 5.1 Variance Due to Nationality Perturbation
Our first research question was to analyse the extent of variation in outputs when nationalities are perturbed in the prompt. For this, we quantify the lexical variance (§ 3.1) in outputs, as measured by word edit distance in Figure 2. We find that model outputs do vary with changing nationalities for both tasks across models. Moreover, these variations are non-trivial and task-dependent, as described below.
Control Experiment: Variance Within Nationalities
In order to ensure that the variance observed across nationalities is non-trivial, i.e. they do not occur because of the non-deterministic nature of generation in models, we also measure the variance within multiple outputs for a particular nationality. We find that the variance for outputs within nationality is consistently lower than the variance across nationalities. We confirm this with ANOVA having a p-value of <0.05 (Appendix A.2).
@@
We find that the nature of the task affects the extent of variation. The median variance for story generation is higher than the median variance for QA for every model. This might be expected as story generation, had longer outputs and being a creative task allows for more adaptations. On the other hand, the difference between the upper and lower quartiles of variance for QA is larger than that for stories. This is likely because QA consists of a wider variety of topics ranging from scientific categories, where limited variations might be expected, to politics and history, that allow more variation in answers than others. For example, answers while explaining ‘elections’ (politics) might vary more as they are operationalized differently across countries, but explaining ‘inertia’ (physics) might not vary as much.
@@ 5.2 Culturally Relevant Words in Outputs
For the topics in the politics category of the QA task, we see words referring to senate houses and political offices of the countries, for example, ‘lok sabha’ and ‘rajya sabha’ in India, ‘bundestag’ for Germany, and ‘meshrano jirga’ and ‘wolesi jirga’ for Afghanistan. The list also features politically polarised issues such as ‘gun’ in America and ‘brexit’ in UK. Another common feature is the names of political parties, such as ‘bjp’ in India, ‘apc’ and ‘pdp’ in Nigeria, and ‘ndp’ in Canada. [^14] Proper nouns here appear lower-cased because we lower-cased all model outputs before calculating the statistics.
Finally, we note that the cultural relevance of all the words on the lists is not obvious (e.g. ‘notably’ in German in Table 2). Moreover, not all topics in the QA setting surface such interpretable lists of culturally relevant words. Especially the lexicon from scientific topics did not reveal interesting examples when inspecting the top-scoring TF-IDF words. This further compliments our earlier finding of output variations being different across tasks.